# An Empirical Study of Kotlin-Java Interactions

### v1

@@ 4.3.2. Java Invoke NPE: fix by clarifying Nullability on Java code
Figure 4 shows commit f240ed6333https://github.com/ant-druha/intellij-powershell/commit/f240ed6 from the intellij-powershell project, which is a good example of the importance of Java nullability annotation when interacting with Kotlin. To “fix exception thrown when plugin is enabled but no powershell is found in the path” as stated in the commit comment, @NotNull annotation was added to the powerShellExecPath Java method (shown in lime color) and Kotlin code which invoked this method was also changed accordingly.
@@ 4.3.5. Kotlin Inherit Java Nullability Inconsistency: fix by ensuring parent-child API consistent
As we can see from the following code, DeclarationSubject class is written in Kotlin and it extends the Subject Java class. The second parameter Declarations in the child Kotlin class is not null while the parent Java class Subject specify the second parameter actual as @Nullable. There exists an inconsistency between these two constructor methods, though this does not cause any problems before version 1.1.3 as the callers of these constructors in the Factory interface does not specify @Nullable in its parameter too.
@@ 5.3. Threats to Validity
One threat is the accuracy of our dependency extraction tool. DependExtractor is based on the Kotlin compiler and dependency types are defined according to previous research. Though we added test cases and manually validate the tool in several small projects, we cannot guarantee that it is 100% accurate for extracted dependencies. We plan to conduct more tests and validations of the tool in the future work.
Another threat is that we only investigated 23 real-world projects with Kotlin as the primary languages. It is possible that additional categories of problematic Kotlin-Java or Java-Kotlin interactions may be discovered in other projects. Exploring this possibility remains part of our future work.
@@ 7. Conclusions
In this paper, we conducted an empirical study of Kotlin-Java interactions in 23 real-world projects. Our results demonstrate that Kotlin and Java interact frequently across the entire systems, confirming the extensive interoperability of Kotlin and Java. However, our study also reveals that source files involved in such cross-language systems have more commits compared to source files involved in only a single language.
Furthermore, we identified seven common mistakes in Kotlin-Java interactions, along with their corresponding fixing strategies. Unlike code smells, which are detected by rules and often have high false positives, these mistakes were extracted from commits in Git repositories and have been shown to require more maintenance effort. We believe that recognizing these identified mistakes in Kotlin-Java interactions can assist developers in effectively improving such cross-langauage interactions.

### v2

@@ 4.3.2. Java Invoke NPE: fix by clarifying Nullability on Java code
Figure 3 shows commit f240ed6333https://github.com/ant-druha/intellij-powershell/commit/f240ed6 from the intellij-powershell project, which is a good example of the importance of Java nullability annotation when interacting with Kotlin. To “fix exception thrown when plugin is enabled but no powershell is found in the path” as stated in the commit comment, @NotNull annotation was added to the powerShellExecPath Java method (shown in lime color) and Kotlin code which invoked this method was changed accordingly.
@@ 4.3.5. Kotlin Inherit Java Nullability Inconsistency: fix by ensuring parent-child API consistent
As we can see from the following code, DeclarationSubject class is written in Kotlin and it extends the Subject Java class. The second parameter Declarations in the child Kotlin class is not null while the parent Java class Subject specify the second parameter actual as @Nullable. There exists an inconsistency between these two constructor methods, though this does not cause any problems before version 1.1.3 as the callers of these constructors in the Factory interface does not specify @Nullable in its parameter.
@@ 5.3. Threats to Validity
The first threat is the accuracy of our dependency extraction tool. DependExtractor is based on the Kotlin compiler and dependency types are defined according to previous research. Though we built the ground truth of a Kotlin-Java project and validated our tool can achieve a relatively high precision and recall, we cannot guarantee that it is 100% accurate. We will continue to validate and improve our tool’s accuracy in the future.
Another threat is that we associated #commit and LOC with file paths. A file involved in K-J dependencies may undergo activities unrelated to fixing interaction problems, or these activities may be mixed in a commit. We acknowledge that #commit and LOC may not capture the exact maintenance cost of a file. However, there is no practical way to split the cost and directly measure the exact amount. We hope that the qualitative analysis detailed in Section 4.3 can partly mitigate this threat.
An external validity is that we only investigated 23 real-world projects with Kotlin as the primary languages. It is possible that additional categories of problematic Kotlin-Java or Java-Kotlin interactions may be discovered in other projects. Our future work will aim to include more software systems and industry examples to further validate our findings.
@@ 7. Conclusions
In this paper, we conducted an empirical study of Kotlin-Java interactions in 23 real-world projects. Our results demonstrate that Kotlin and Java interact frequently, confirming the extensive interoperability of Kotlin and Java. However, our study also reveals that source files involved in such cross-language systems have more commits compared to those involved in only a single language.
Furthermore, we identified seven common mistakes in Kotlin-Java interactions, along with their corresponding fixing strategies. Unlike code smells, which are detected by rules and often have high false positives, these mistakes were extracted from commits in Git repositories and have been shown to require more maintenance effort. We believe that recognizing these identified mistakes in Kotlin-Java interactions can assist developers in effectively improving such cross-language interactions.
Our future work will focus on the following: first, investigating more systems and industry examples to further validate our findings; second, extending this empirical study to an automatic tool that can detect and fix problematic Kotlin-Java interactions; and third, enhancing our DependExtractor tool to support more cross-language dependencies.